OpenStudy (anonymous):

What is the oblique asymptote of the function f(x) = x^2+5x+6/x+4

OpenStudy (anonymous):

y=x+9, y=x-9, y=x+1,y=x-1

OpenStudy (tkhunny):

1) The expression you have written has no aoblque asymptote. It has onyl a vertical asymptote at x = 0. 2) If you meant to write this, \(\dfrac{x^{2}+5x+6}{x+4}\), you needed more parentheses. The Olbique Asymptote is most easily found by long division.'\(\dfrac{x^{2}+5x+6}{x+4} = x + 1 + \dfrac{2}{x+4}\). It is now staring at you.